# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

### v1

@@ II.2 Helical AB interferometer
Figure 4: The scattering amplitudes entering scattering matrix of contacts, Eq. (3), which are modelled as a single-channel spinfull wire. It is assumed that there is no spin flip at the contact and the two spin polarizations at the contact (shown by red and blue color) are completely separated. The $t\to 1$ case corresponds to a tunnel contact, and the $t\to 0$ case models a metal contact.
@@
$\hat{S}^{\text{HI}}=\begin{pmatrix}-t&r&0&0\\ r&t&0&0\\ 0&0&-t&r\\ 0&0&r&t\\ \end{pmatrix}\,,\quad t^{2}+r^{2}=1,$ (3)
@@
Importantly, due to the topological protection, backscattering appears in the HI only in presence of magnetic defects or magnetization of the leads. We assume in this paper that the leads are non-magnetic but there is a static magnetic defect in one of the arms of the interferometer. A comment is in order here. As we already mentioned, much attention was drawn recently to the infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. For the isotropic exchange interaction between the impurity and the HES, the impurity’s magnetic moment relaxes within the so-called Korringa time towards the local direction of the electron spin in the HES and, as a consequence, the interaction between the impurity and the HES becomes completely ineffective (see discussion in Väyrynen and Glazman (2017)). Accordingly, the problem of the noise intensity at zero frequency in the HES with a single impurity makes sense only in the presence of an anisotropic exchange interaction Kurilovich et al. (2019a) (or an external magnetic field acting on a dynamic impurity Probst et al. (2022)), while in the limit of isotropic exchange interaction Fano factor is singular Kurilovich et al. (2019a), i.e. the result for the $\mathcal{F}$ depends on the order in which the constants responsible for the anisotropy tend to zero (see also discussion in Nagaev et al. (2018))
@@
Such a defect, considered below, ensures the existence of a magnetic field in a small region of the HES, i.e. allows elastic backward scattering without tunneling coupling between the HES and defect.
Specifically, we assume the presence of magnetic defect (MD) at the edge of the TI, which can be, for example, a potential dielectric ferromagnetic point contact with high magnetic stiffness and with the magnetic moment, whose direction is determined by uniaxial anisotropy and the demagnetization tensor of the ferromagnet. We describe the electron scattering on such MD by the scattering matrix
$\hat{S}_{\text{MD}}=\begin{pmatrix}\cos\theta e^{i\alpha}&i\sin\theta~{}e^{i\varphi}\\ i\sin\theta~{}e^{-i\varphi}&\cos\theta e^{-i\alpha}\\ \end{pmatrix}\,,$ (5)
that allows for backward scattering, and neglect the back influence of the HES on the parameters of $S_{M}.$ The backward scattering rate, $R_{\theta}=\sin^{2}\theta,$ is determined by the quantity $\theta$ while the phase $\varphi$ has the meaning of the backward scattering phase on the MD.
Here we derive analytical expressions describing anticrossing of quantum levels of right- and left-moving electrons on the example of a single impurity placed in the upper shoulder. We consider interferometer with the lengths of the upper and lower shoulders given by $L_{1}$ and $L_{2},$ respectively. The magnetic impurity is placed at position $x_{0}$ such that $0<x_{0}<L_{1}.$ Using expression for scattering matrix (5), one can easily find transfer matrix of impurity
$\hat{T}_{\text{MD}}=\frac{e^{-i\alpha}}{\cos\theta}\left(\begin{array}[]{cc}1&i\sin\theta e^{-i\xi}\\ -i\sin\theta e^{i\xi}&1\end{array}\right),$ (6)
@@
Note that the possibility of creating static magnetic contacts to the HES has also been discussed in another context Khomitsky et al. (2022). In principle, the interference effects we are interested in could also be observed for a point-like non-magnetic scatterer, taking into account the electron-electron interactionSablikov and Sukhanov (2021). In the latter case, however, an analysis of inelastic effects is required, which is beyond the scope of the current work.
@@ III.1 General equations
In this Section, we calculate $\mathcal{F}$ for setups shown in Fig. 1 assuming that a fixed bias voltage $V$ is applied to leads. We will consider the most interesting and easily realized case:
@@
The current noise is related with fluctuations of the electric current with respect to its average value $\delta\hat{I}(t)=\hat{I}(t)-\langle\hat{I}\rangle$. Here $\hat{I}$ is the current operator (an analytical expression for $\hat{I}$ is given in de Jong and Beenakker (1997); Blanter and Büttiker (2000)). The current correlation function associated with the noise is defined by:
${\mathcal{S}}\left(t-t^{\prime}\right)=\tfrac{1}{2}\langle\delta\hat{I}(t)\delta\hat{I}(t^{\prime})+\delta\hat{I}(t^{\prime})\delta\hat{I}(t)\rangle\,.$
[…]
@@
[…]
where $G_{0}=2e^{2}/h$ conductance quantum and
@@
[…]
@@
$\mathcal{F}=\frac{\text{Tr}\langle\hat{\mathcal{T}}(1-\hat{\mathcal{T}})\rangle_{\varepsilon}}{\text{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}}\,,$ (12)
[…]
@@
${\mathcal{T}}=\text{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}/2.$ (13)
@@
${\mathcal{T}}_{2}=\text{Tr}\langle\hat{\mathcal{T}}\hat{\mathcal{T}}\rangle_{\varepsilon},$ (14)
@@
In the next two subsections we will discuss calculations that allows one to find $\mathcal{F},$ using Eq. (15). Technically, the key idea is to present the matrix of transmission amplitudes in a special form, which allows to perform the energy averaging analytically.
@@ III.2 Conventional AB interferometer
We start with calculation of $\mathcal{F}$ for CI having two arms with the lengths $L/2\pm a$ and two contacts (not necessarily identical) described by scattering matrices with real amplitudes [see Eq.  (2)]. We first find convenient expression for matrix of transmission amplitude (which is a scalar for spinless CI), substitute it into the formula (12), and average over the energy.
@@ III.2.1 Method of calculation
We use the method developed in Ref. Shmakov et al. (2013) to sum amplitudes with different winding numbers. We classify all amplitudes of electrons going out of the ring by their direction, either clockwise, (+)(+)( + ), or anticlockwise, ()(-)( - ). For CI with identical contacts, the amplitudes describing processes of passing the ring in opposite directions without extra revolutions are given by
$\begin{pmatrix}\beta_{0}^{+}\\ \beta_{0}^{-}\end{pmatrix}=t_{\text{in}}t_{\text{out}}\begin{pmatrix}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{pmatrix}\,.$ (16)
All other trajectories finishing at the right lead can be obtain recursively, multiplying by a matrix describing processes which increase number of revolutions, clockwise or counterclockwise, by unity (for a more detailed discussion, see Ref. Shmakov et al. (2013)):
$\begin{pmatrix}\beta_{n+1}^{+}\\ \beta_{n+1}^{-}\end{pmatrix}=\hat{A}\begin{pmatrix}\beta_{n}^{+}\\ \beta_{n}^{-}\end{pmatrix},$
$\hat{A}=e^{ikL}\begin{pmatrix}t^{2}e^{-i2\pi\phi}+t_{b}^{2}e^{i2ka}&&tt_{b}(e^{-i2\pi\phi}+e^{i2ka})\\ tt_{b}(e^{i2\pi\phi}+e^{-i2ka})&&t^{2}e^{i2\pi\phi}+t_{b}^{2}e^{-i2ka}\end{pmatrix}.$
@@
${t}^{\rm CI}(\varepsilon)=\sum_{n=0}^{\infty}\alpha\hat{A}^{n}\beta_{0}=\alpha(1-\hat{A})^{-1}\beta_{0}\,,\quad\alpha=\left(\begin{matrix}1\\ 1\end{matrix}\right).$ (17)
Having particular expressions for $t^{\rm CI}(\varepsilon),$ we get ${\cal T}^{\text{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{2}$ and ${\cal T}_{2}^{\text{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{4}$. The averaging over $\varepsilon$ is reduced to integration over the unit circle in complex plane of $z=e^{ikL}$ and is easily performed by residues, see App A. Most interesting and physically transparent limiting cases are presented below. In the most general case of interferometer with contacts of different strength, the obtained analytical expressions are rather cumbersome and we present them in the Appendix B.
@@ III.2.2 Results (symmetric setup, identical contacts)
Figure 5: The dependence of Fano factor, $\mathcal{F}^{\text{CI}}(\phi,t_{\text{b}})$, on the amplitude of backscattering on the contact, $t_{\text{b}}$, and the magnetic flux, $\phi$, as described by Eqs. (2), (18), (19).
Figure 6: The transmission coefficient (blue curves) and the Fano factor (red curves) for conventional interferometer with two identical contacts. In symmetric interfererometer with $a=0$ and weak tunneling coupling ($\gamma=0.1$) there is antiresonance in $\mathcal{T}^{\rm CI}$ and resonance in $\mathcal{F}^{\rm CI}$ at half-integer values of the flux [panels (a) and (b)]. In asymmetric interferometer with $a\neq 0$ and weak tunneling coupling ($\gamma=0.1$, $k_{F}a=1.0$), there also appear peaks at integer values of the flux [panels (c) and (d)]. With increasing $a,$ peaks at $\phi=1/2+n$ decrease in amplitude and become narrower, while peaks at $\phi=n$ grow and become wider (here, n is integer).
@@
${\mathcal{T}}^{\text{CI}}(\phi,\gamma)=\frac{2\gamma\cos^{2}(\pi\phi)}{\gamma^{2}+\cos^{2}(\pi\phi)}\,.$ (18)
Using formulas from Appendix A we find the Fano factor in the form
$\mathcal{F}^{\text{CI}}(\phi,\gamma)\!=1\!-\!\frac{\cos^{2}(\pi\phi)\!\left[5\gamma^{2}\!+\!\gamma^{4}\!+\!(1\!+\!\gamma^{2})\cos^{2}(\pi\phi)\right]}{2\left[\gamma^{2}+\cos^{2}(\pi\phi)\right]^{2}}.$ (19)
Equation (19) is illustrated in in Fig. 5, where the dependence of Fano factor on the magnetic flux and on the backscattering on the contact, $|t_{\text{b}}|=\gamma/(1+\gamma).$ For three types of tunneling contacts shown in Fig. 3, Eq. (19) simplifies:
$\mathcal{F}^{\text{CI}}(\phi,0)=$ $\tfrac{1}{2}\,,$ (20)
$\mathcal{F}^{\text{CI}}(\phi,\infty)=$ $1-\tfrac{1}{2}\cos^{2}(\pi\phi)\,,$
$\mathcal{F}^{\text{CI}}(\phi,1)=$ $\frac{4\sin^{2}(\pi\phi)}{(\cos(2\pi\phi)+3)^{2}}\,.$
The value $1/2$ for $\gamma\to 0$ is a standard value for the Fano factor of tunneling level. We also note that the “metallic” contacts with $\gamma=1$, shows both minimal value $\mathcal{F}^{\text{CI}}=0$, at $\phi=0$, and to the maximal value, $\mathcal{F}^{\text{CI}}=1$, at $\phi\to 1/2.$
The dependencies of ${\mathcal{T}}^{\text{CI}}$ and $\mathcal{F}^{\text{CI}}$ on the magnetic flux are shown in Fig. 6a, 6b. The latter figure yieled cross-section of Fig. 5 for fixed $\gamma.$ We see that the transmission coefficient and the Fano factor have, respectively, sharp antiresonance and sharp resonance at $\phi=1/2$ (and, consequently, at all half-integer values of the flux).
Figure 7: Parametric plot $(\mathcal{F}^{\rm CI},\mathcal{T}^{\rm CI})$ for fixed $\phi$ and $0<\gamma<\infty$ (a), for fixed $|t_{\rm b}|=\gamma/(1+\gamma)$ and $10<\phi<$ 1 (b).
Several interesting properties of $\mathcal{F}^{\text{CI}}(\phi,\gamma)$ are to be mentioned here. Firstly,
$\mathcal{F}^{\text{CI}}(0,\gamma)=\mathcal{F}^{\text{CI}}(0,1/\gamma)=\frac{1}{2}\left(\frac{1-\gamma^{2}}{1+\gamma^{2}}\right)^{2},$ (21)
@@
$\mathcal{T}^{\text{CI}}(0,\gamma)=\mathcal{T}^{\text{CI}}(0,1/\gamma)=\frac{2\gamma}{1+\gamma^{2}}.$
@@
$\mathcal{F}^{\text{CI}}=\tfrac{1}{2}\left[1-(\mathcal{T}^{\text{CI}})^{2}\right]\,.$ (22)
This equation is shown by red curve in Fig. 7 and is valid both for $\gamma<1$ (i.e. for contacts shown in Fig. 3a) and for $\gamma>1$ (i.e. for contacts shown in Fig. 3b). Specifically, changing $\gamma$ from $\gamma=0$ to $\gamma=\infty,$ we pass the red curve twice: once in the interval $10<\gamma<$ 1 and once again for $1<\gamma<\infty.$ This is clearly seen from the cross-section of Fig. at $\phi=0.$ Having in mind to compare below with the HI, we note that the role of the amplitude $t_{\text{b}}$ is somewhat similar to the scattering on the magnetic impurity in case of HI, considered below. For the case of CI, dependence on $|t_{\rm b}|$ at $\phi=0$ (i.e. on the red curve Fig. 3a) directly follows from Eq. (21):
@@
$\mathcal{T}^{\text{CI}}(1/2,\gamma)=0,\quad\mathcal{F}^{\text{CI}}(1/2,\gamma)=1.$ (23)

### v2

@@ II.2 Helical AB interferometer
Figure 4: The scattering amplitudes entering scattering matrix of contacts, Eq. (4), which are modeled as a single-channel spinful wire. It is assumed that there is no spin flip at the contact and the two spin polarizations at the contact (shown by red and blue color) are completely separated. The $t\to 1$ case corresponds to a tunnel contact, and the $t\to 0$ case models a metal contact. Contact region with area $S_{0}$ is shown by gray color. We assume that $BS_{0}\ll\Phi_{0},$ so that the field does not affect S-matrix.
@@
$\hat{S}^{\mathrm{HI}}=\left(\begin{array}[]{cccc}-t&r&0&0\\ r&t&0&0\\ 0&0&-t&r\\ 0&0&r&t\\ \end{array}\right)\,,\quad t^{2}+r^{2}=1,$ (4)
@@
Due to the topological protection, backscattering appears in the HI only in presence of a defect that violate time-reversal symmetry, in particular, in presence of magnetic defects or in case of magnetization of the leads. We assume in this paper that the leads are non-magnetic, but there is a static backscattering defect in one of the arms of the interferometer. As we already mentioned, much attention was drawn recently to the infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. For the isotropic exchange interaction between the impurity and the HES, the impurity’s magnetic moment relaxes within the so-called Korringa time towards the local direction of the electron spin in the HES and, as a consequence, the interaction between the impurity and the HES becomes completely ineffective (see discussion in Väyrynen and Glazman (2017)). Accordingly, the problem of the noise intensity at zero frequency in the HES with a single impurity makes sense only in the presence of an anisotropic exchange interaction Kurilovich et al. (2019a) (or an external magnetic field acting on a dynamic impurity Probst et al. (2022)), while in the limit of isotropic exchange interaction Fano factor is singular Kurilovich et al. (2019a), i.e. the result for the $\mathcal{F}$ depends on the order in which the constants responsible for the anisotropy tend to zero (see also discussion in Nagaev et al. (2018))
@@
Such a defect ensures the existence of a magnetic field in a small region of the HES, i.e. it allows elastic backward scattering without tunneling coupling between the HES and defect. Possible experimental realization is, for example, a potential dielectric ferromagnetic point contact with high magnetic stiffness and with the magnetic moment, whose direction is determined by uniaxial anisotropy and the demagnetization tensor of the ferromagnet. The possibility of creating static magnetic contacts to the HES has also been discussed Khomitsky et al. (2022).
Importantly, backscattering can occur not only due to magnetic defect but also due to the interaction with a charged puddle in the bulk of the topological insulator Väyrynen et al. (2013) or due to a point-like non-magnetic scatterer, with account for the electron-electron interaction Sablikov and Sukhanov (2021).
Here, we assume the presence of backscattering static defect (BD) at the edge of the TI, which is described by the scattering matrix of the most general type
$\hat{S}_{\mathrm{BD}}=\left(\begin{array}[]{cc}\cos\theta e^{i\alpha}&i\sin\theta~e^{i\varphi}\\ i\sin\theta~e^{-i\varphi}&\cos\theta e^{-i\alpha}\\ \end{array}\right)\,.$ (6)
For the case of magnetic defect, we neglect the back influence of the HES on the parameters of $S_{\mathrm{BD}}$. The backward scattering rate, $R_{\theta}=\sin^{2}\theta,$ is determined by the quantity $\theta$ while the phase $\varphi$ has the meaning of the backward scattering phase on the BD.
Here we derive analytical expressions for a single BD placed in the upper shoulder. We consider interferometer with the lengths of the upper and lower shoulders given by $L_{1}$ and $L_{2},$ respectively. The BD is placed at position $x_{0}$ such that $0<x_{0}<L_{1}.$ Using expression for scattering matrix (6), one can easily find transfer matrix of the defect
$\hat{T}_{\mathrm{BD}}=\frac{e^{-i\alpha}}{\cos\theta}\left(\begin{array}[]{cc}1&i\sin\theta e^{-i\xi}\\ -i\sin\theta e^{i\xi}&1\end{array}\right),$ (7)
@@ III.1 General equations
In this Section, we calculate $\mathcal{F}$ for setups shown in Fig. 1 assuming that a fixed bias voltage $V$ is applied to leads. (In this paper, we assume that the external impedance is zero and therefore the voltage does not fluctuate. For a finite impedance both current and voltage fluctuate as discussed in detail in Ref. Blanter and Büttiker (2000).)
We will consider the most interesting and easily realized case:
@@
The current noise is related with fluctuations of the electric current with respect to its average value $\delta\hat{I}(t)=\hat{I}(t)-\langle\hat{I}\rangle$. Here $\hat{I}$ is the current operator (an analytical expression for $\hat{I}$ is given in de Jong and Beenakker (1997); Blanter and Büttiker (2000)).
The current correlation function associated with the noise is defined by:
${\mathcal{S}}\left(t-t^{\prime}\right)=\frac{1}{2}\langle\delta\hat{I}(t)\delta\hat{I}(t^{\prime})+\delta\hat{I}(t^{\prime})\delta\hat{I}(t)\rangle\,.$
The Fourier transform of $\mathcal{S}$} gives an expression for the noise power: $S(\omega)=2\int_{-\infty}^{\infty}dt\,e^{i\omega t}{S(t)}$ [the factor $2$ in this expression is a matter of convention, see Eq. (1) in Ref. de Jong and Beenakker (1997) and the comment after Eq. (49) in Ref. Blanter and Büttiker (2000)].
@@
$\mathcal{S}(\omega=0)$ $=2G_{0}$ (9)
$\times\int_{\mu}^{\mu+eV}d\varepsilon\,{\rm Tr}\left[{}\hat{\mathcal{T}}\left(\varepsilon\right)\left(1-\hat{\mathcal{T}}\left(\varepsilon\right)\right)\right]\,$,
where $G_{0}=e^{2}/h$ conductance quantum and
@@
$eI=G_{0}\int_{\mu}^{\mu+eV}d\varepsilon\,{\rm Tr}\left[\hat{\mathcal{T}}\left(\varepsilon\right)\right]\,$, (11)
$\mathcal{F}=\frac{\mathcal{S}(\omega=0)}{2eI}=\frac{\int_{\mu}^{\mu+eV}d\varepsilon\,\mathrm{Tr}\left[\hat{\mathcal{T}}(1-\hat{\mathcal{T}})\right]}{\int_{\mu}^{\mu+eV}d\varepsilon\,\mathrm{Tr}[\hat{\mathcal{T}}]}\,$. (12)
@@
$\mathcal{F}=\frac{\mathrm{Tr}\langle\hat{\mathcal{T}}(1-\hat{\mathcal{T}})\rangle_{\varepsilon}}{\mathrm{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}}\,$, (13)
where the averaging is taken over a temperature window in the vicinity of the Fermi level in the limit $T\gg\Delta$. In what follows we assume the linearized form of the spectrum, with $\varepsilon=v_{\rm F}k$, so that the energy averaging is reduced to calculation of the integral $\left\langle\cdots\right\rangle_{\varepsilon}=\Delta^{-1}\int_{0}^{\Delta}d\varepsilon\,(\cdots)=\frac{L}{2\pi}\int_{0}^{2\pi/L}dk\,(\cdots)$.
@@
${\mathcal{T}}=\mathrm{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}/2.$ (14)
@@
${\mathcal{T}}_{2}=\mathrm{Tr}\langle\hat{\mathcal{T}}\hat{\mathcal{T}}\rangle_{\varepsilon},$ (15)
@@
In the next two subsections we will discuss calculations that allow one to find $\mathcal{F},$ using Eq. (16). Technically, the key idea is to present the matrix of transmission amplitudes in a special form, which allows performing the energy averaging analytically.
@@ III.2 Conventional AB interferometer
We start with calculation of $\mathcal{F}$ for CI having two arms with the lengths $L/2\pm a$ and two contacts (not necessarily identical) described by scattering matrices with real amplitudes [see Eq.  (2)]. We first find convenient expression for the matrix of transmission amplitude (that consists of one element for spinless CI, also conductance quantum is $G_{0}=e^{2}/h$ in this case), substitute it into the formula (13), and average over the energy.
@@ III.2.1 Method of calculation
We use the method developed in Ref. Shmakov et al. (2013) to sum amplitudes with different winding numbers. We classify all amplitudes of electrons going out of the ring by their direction, either clockwise, $(+)$, or anticlockwise, ()(-). For CI with identical contacts, the amplitudes describing processes of passing the ring in opposite directions without extra revolutions are given by
$\left(\begin{array}[]{c}\beta_{0}^{+}\\ \beta_{0}^{-}\end{array}\right)=t_{\mathrm{in}}t_{\mathrm{out}}\left(\begin{array}[]{c}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{array}\right)\,.$ (17)
All other trajectories finishing at the right lead can be obtained recursively, multiplying by a matrix describing processes which increase number of revolutions, clockwise or counterclockwise, by unity (for a more detailed discussion, see Ref. Shmakov et al. (2013)):
$\left(\begin{array}[]{c}\beta_{n+1}^{+}\\ \beta_{n+1}^{-}\end{array}\right)=\hat{A}\left(\begin{array}[]{c}\beta_{n}^{+}\\ \beta_{n}^{-}\end{array}\right),$ (18)
$\hat{A}=e^{ikL}$
$\times\left(\begin{array}[]{cc}t^{2}e^{-i2\pi\phi}+t_{b}^{2}e^{i2ka}&tt_{b}(e^{-i2\pi\phi}+e^{i2ka})\\ tt_{b}(e^{i2\pi\phi}+e^{-i2ka})&t^{2}e^{i2\pi\phi}+t_{b}^{2}e^{-i2ka}\end{array}\right).$
@@
${t}^{\rm CI}(\varepsilon)=\sum_{n=0}^{\infty}\alpha\hat{A}^{n}\beta_{0}=\alpha(1-\hat{A})^{-1}\beta_{0}\,$, (19)
$\alpha=\left(\begin{array}[]{c}1\\ 1\end{array}\right).$
Having particular expressions for $t^{\rm CI}(\varepsilon),$ we get ${\cal T}^{\mathrm{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{2}$ and ${\cal T}_{2}^{\mathrm{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{4}$. The averaging over $\varepsilon$ is reduced to integration over the unit circle in complex plane of $z=e^{ikL}$ and is easily performed by residues, see A. Most interesting and physically transparent limiting cases are presented below. In the most general case of interferometer with contacts of different strength, the obtained analytical expressions are rather cumbersome and we present them in the B.
@@ III.2.2 Results (symmetric setup, identical contacts)
Figure 5: The dependence of Fano factor, $\mathcal{F}^{\mathrm{CI}}(\phi,t_{\mathrm{b}})$, on the amplitude of backscattering on the contact, $t_{\mathrm{b}}$, and the magnetic flux, $\phi$, as described by Eqs. (2), (20), (21).
Figure 6: The transmission coefficient (blue curves) and the Fano factor (red curves) for conventional interferometer with two identical contacts. In symmetric interferometer with $a=0$ and weak tunneling coupling ($\gamma=0.1$) there is antiresonance in $\mathcal{T}^{\rm CI}$ and resonance in $\mathcal{F}^{\rm CI}$ at half-integer values of the flux [panels (a) and (b)]. In asymmetric interferometer with $a\neq 0$ and weak tunneling coupling ($\gamma=0.1$, $k_{F}a=1.0$), there also appear peaks at integer values of the flux [panels (c) and (d)]. With increasing $a,$ peaks at $\phi=1/2+n$ decrease in amplitude and become narrower, while peaks at $\phi=n$ grow and become wider (here, n is integer). Shape of the curve $\mathcal{T}^{\rm CI}(\phi)$ strongly depends on the asymmetry of the contacts.
@@
${\mathcal{T}}^{\mathrm{CI}}(\phi,\gamma)=\frac{2\gamma\cos^{2}(\pi\phi)}{\gamma^{2}+\cos^{2}(\pi\phi)}\,.$ (20)
Using formulas from A we find the Fano factor in the form (this equation was presented in Ref. Niyazov et al. (2024) without derivation)
$\mathcal{F}^{\mathrm{CI}}(\phi,\gamma)\!=1\!$ (21)
$-\!\frac{\cos^{2}(\pi\phi)\!\left[5\gamma^{2}\!+\!\gamma^{4}\!+\!(1\!+\!\gamma^{2})\cos^{2}(\pi\phi)\right]}{2\left[\gamma^{2}+\cos^{2}(\pi\phi)\right]^{2}}.$
Equation (21) is illustrated in Fig. 5, where the dependence of Fano factor on the magnetic flux and on the backscattering on the contact, $|t_{\mathrm{b}}|=\gamma/(1+\gamma).$ For three types of tunneling contacts shown in Fig. 3, Eq. (21) simplifies:
$\mathcal{F}^{\mathrm{CI}}(\phi,0)=$ $\frac{1}{2}\,$, (22)
$\mathcal{F}^{\mathrm{CI}}(\phi,\infty)=$ $1-\frac{1}{2}\cos^{2}(\pi\phi)\,$,
$\mathcal{F}^{\mathrm{CI}}(\phi,1)=$ $\frac{4\sin^{2}(\pi\phi)}{(\cos(2\pi\phi)+3)^{2}}\,.$
The value $1/2$ for $\gamma\to 0$ is a standard value for the Fano factor of tunneling level. We also note that the “metallic” contacts with $\gamma=1$, shows both minimal value $\mathcal{F}^{\mathrm{CI}}=0$, at $\phi=0$, and to the maximal value, $\mathcal{F}^{\mathrm{CI}}=1$, at $\phi\to 1/2.$
The dependencies of ${\mathcal{T}}^{\mathrm{CI}}$ and $\mathcal{F}^{\mathrm{CI}}$ on the magnetic flux are shown in Fig. 6a, 6b. The latter figure yields cross-section of Fig. 5 for fixed $\gamma.$ In Fig. 6a we see that in the symmetrical interferometer, $a=0,$ the transmission coefficient and the Fano factor have, respectively, sharp antiresonance and sharp resonance at $\phi=1/2$ (and, hence, at any half-integer flux value).
Figure 7: Parametric plot $(\mathcal{F}^{\rm CI},\mathcal{T}^{\rm CI})$ for fixed $\phi$ and $0<\gamma<\infty$ (a), for fixed $|t_{\rm b}|=\gamma/(1+\gamma)$ and $0<\phi<1$ (b).
Several interesting properties of $\mathcal{F}^{\mathrm{CI}}(\phi,\gamma)$ are to be mentioned here. Firstly,
$\mathcal{F}^{\mathrm{CI}}(0,\gamma)=\mathcal{F}^{\mathrm{CI}}(0,1/\gamma)=\frac{1}{2}\left(\frac{1-\gamma^{2}}{1+\gamma^{2}}\right)^{2},$ (23)
@@
$\mathcal{T}^{\mathrm{CI}}(0,\gamma)=\mathcal{T}^{\mathrm{CI}}(0,1/\gamma)=\frac{2\gamma}{1+\gamma^{2}}.$
@@
$\mathcal{F}^{\mathrm{CI}}=\frac{1}{2}\left[1-(\mathcal{T}^{\mathrm{CI}})^{2}\right]\,.$ (24)
This equation is shown by red curve in Fig. 7 and is valid both for $\gamma<1$ (i.e. for contacts shown in Fig. 3a) and for $\gamma>1$ (i.e. for contacts shown in Fig. 3b). Specifically, changing $\gamma$ from $\gamma=0$ to $\gamma=\infty,$ we pass the red curve twice: once in the interval $0<\gamma<1$ and once again for $1<\gamma<\infty.$ This is clearly seen from the cross-section of Fig. 5 at $\phi=0.$ Having in mind to compare below with the HI, we note that the role of the amplitude $t_{\mathrm{b}}$ is somewhat similar to the scattering on the magnetic impurity in case of HI, considered below. For the case of CI, dependence on $|t_{\rm b}|$ at $\phi=0$ (i.e. on the red curve Fig. 3a) directly follows from Eq. (23):
@@
$\mathcal{T}^{\mathrm{CI}}(1/2,\gamma)=0,\quad\mathcal{F}^{\mathrm{CI}}(1/2,\gamma)=1.$ (25)